### v1

# Loss of Control Prevention of an Agile Aircraft: Dynamic Command Saturation Approach

Loss of Control Prevention of an Agile Aircraft: Dynamic Command Saturation Approach
Ege Ç. Altunkaya  [^1] Research Assistant, Aviation Institute, Aerospace Research Center, Istanbul Technical University, Istanbul, Türkiye, 34469. and Akın Çatak  [^2] Research Assistant, Aviation Institute, Aerospace Research Center, Istanbul Technical University, Istanbul, Türkiye, 34469. Aerospace Research Center, Istanbul Technical University, Istanbul, Türkiye, 34469 Emre Koyuncu  [^3] Assoc. Prof., Aerospace Research Center, Istanbul Technical University, Istanbul, Türkiye, 34469. Aerospace Research Center, Istanbul Technical University, Istanbul, Türkiye, 34469 İbrahim Özkol  [^4] Prof., Aerospace Research Center, Istanbul Technical University, Istanbul, Türkiye, 34469. Aerospace Research Center, Istanbul Technical University, Istanbul, Türkiye, 34469
1 Introduction
[…]
The loss of control, a leading factor in fatal crashes, poses a critical threat to flight safety by jeopardizing the flight stability [1]. Numerous factors contribute to this loss, ranging from adverse onboard conditions to unexpected vehicle upsets [2, 3]. The diverse array of reasons and their interactions inherently elevate the risk of control loss, as extensively outlined in terms of fatal crash statistics linked to the control loss in [2] and [3]. Moreover, a detailed analysis of the causes behind the control loss reveals that the pilot-induced loss of control, a significant contributing factor, involves abrupt and aggressive maneuvers that can trigger a destabilized flight [3]. Despite the desirability of performing aggressive maneuvers without compromising safety, it remains constrained by the aircraft’s aerodynamic characteristics, controllability, and its flight dynamics. Additionally, according to [4], reasons for loss of control can be categorized under five major headlines: stall, sideslip-induced rolls, rolls from other causes, pilot-induced oscillation, and yaw.
Thereby designing an aircraft with a high-performance flight control law is not always sufficient and does not imply that the aircraft can perform any sort of maneuver. There are numerous restrictions that hinder an agile aircraft from conducting agile maneuvers, such as high angle of attack kinematics, inertia-coupling, and flight and/or dynamic envelope violation. The preceding factors are so hazardous that they may trigger an upset condition such as loss of control, departure, spin, stall, or even structural destruction. In summary, loss of control refers to a flight dynamics and control issue where the aircraft’s motion deviates from expected operating conditions, and this problem involves non-linear effects like nonlinear kinematic and inertial coupling, which cause enormous responses to small changes in state variables or result in the oscillatory or divergent behavior. As a consequence, it often leads to significant angular rates and displacements, making it challenging to maintain the desired heading, altitude, and level flight [5]. Therefore, to prevent any upset condition, one has to delve into presumable reasons for these catastrophic consequences.
Since loss of control still persists as a critical issue, numerous studies inherently exist to address it from various perspectives. Hereby, ensuring safe flight has been extensively studied through flight envelope protection [6, 7, 8, 9, 10, 11, 12, 13, 14], prevention of loss of control [15, 9, 16], and implementation of command/reference governor methods [17, 18]. Lombaerts et al. [9] detail the design process of the adaptive algorithm and its integration into a piloted simulator for evaluation. The paper discusses the methodology employed in simulator testing, including the scenarios used to assess the algorithm’s performance. The boundaries of the flight envelope are determined adaptively, and the flight envelope-aware flight control system serves to prevent loss of control. Finally, they implement the proposed architecture into a simulator environment, belonging to a commercial airline aircraft, for further examination. Yu et al. [13] present an approach for designing safety control systems for aircraft that incorporate flight envelope protection and reference command generation. The approach is based on a nonlinear control framework that combines feedback control with feedforward reference commands to achieve robust and safe flight performance under a variety of operating conditions. The results show that the approach is effective at maintaining stable flight and protecting the aircraft from operating outside of its safe flight envelope. Despite the existence of adaptive flight envelope determination methods, there also exist other online safe flight envelope estimation studies [19, 20] besides the ANN-based flight envelope estimation approach [21].
Nevertheless, various studies strongly emphasize the distinction between the flight envelope and dynamic envelope of an aircraft [22, 1]. In other words, the dynamic envelope defines the region within the aircraft’s state space where safe controllability is assured, preventing the occurrence of events leading to loss of control. Consequently, the notion of ensuring flight conditions solely within the traditional flight envelope for the prevention of the loss of control becomes a matter of debate. Particularly, Zheng et al. [22] study the determination of the dynamic envelope of the baseline aircraft, NASA’s Generic Transport Model. They generate the dynamic envelope using differential manifold theory and claim that this method is able to capture the coupling effects more accurately. Moreover, a correlation between the dynamic envelope and control coefficient or pilot command is established. Finally, based on the dynamic envelope enlargement strategy, an envelope protection architecture is proposed and assessed.
On the other hand, various studies aim to estimate the maneuvering envelope through reachability analysis, such as [23, 24, 15, 20]. However, these approaches are not suitable for the real-time adaptation due to the high computational cost associated with numerically solving the Hamilton-Jacobi-Bellman equations. Furthermore, in [1, 25], loss of control is thoroughly investigated in the sense of nonlinear dynamics, and bifurcation analysis is put forth. However, of course, bifurcation analyses can only be done in the design phase of the aircraft.
Last but not least, maneuver limiting is also addressed in various studies using the command and reference governor approaches. Simon et al. [17] propose a method for limiting the angle of attack and load factor using the command and reference governors. The implementation has been carried out in the simulation environment used for the Jas39 Gripen fighter aircraft. As a result of the study, a remarkable degradation in overshoot is observed in the angle of attack and load factor responses at their limit values, constrained by the corresponding flight envelope. A similar effort has been undertaken by [18] for the flight envelope protection during high-angle-of-attack maneuvers. Albeit the proposed methods are successful in achieving their primary objectives, i.e., flight envelope protection, a priori flight envelope information is a must; otherwise, command and reference governors cannot determine where to stop the aircraft prior to entering an upset condition.
From an industrial perspective, the primary goal is "care-free" maneuvering without compromising performance and safety, as emphasized in [25]. However, the certification remains a constant requirement in the industry, making conventional methods the preferred choice for safety. Traditional state limiters, such as angle of attack, load factor, bank angle, pitch angle, and roll rate limiters, are still in use. These limiters are scheduled based on the flight-critical parameters like Mach number, dynamic pressure, angle of attack, and/or roll rate. Nevertheless, designing these state limiters for safe flight requires a labor-intensive effort. Extensive aerodynamic analyses are necessary to investigate the stability characteristics of the designed aircraft. For this purpose, several helpful notions (not limited to these), such as lateral control departure parameter and $C_{{n_{\beta}}_{dyn}}$, are adopted. Careful analysis of these parameters is of significant importance to determine the departure angle of attack value, and their variations with changing flight circumstances must be considered for scheduling. Importantly, these evaluations are not limited to the early design phases but must be extended and verified through the flight tests. Consequently, these aspects contribute to the labor and cost excessively.
According to the inferences drawn from the previous research, as elaborated previously, the issue of the loss of control has been discussed within the context of the flight envelope protection. Additionally, a priori flight envelope databases generally exist in these studies. Consequently, in such investigations, the aircraft is pre-aware of where to avoid based on the pre-known flight envelope. However, confining the concept of the loss of control solely within the flight envelope protection does not seem inclusive, and the consideration of extreme and coupling maneuvers has been lacking in most previous research. Based on these evaluations, the followings are the contributions of this study:
  • A priori information about where to avoid from does not exist in this study; however, this information is derived in an online manner by merging the Lyapunov’s direct method and inverse dynamics of the aircraft.
  • Controllability investigation has never been handled using the incremental attainable moment set concept. Both the derivation of the incremental attainable moment set and its utilization for an online controllability assessment are discussed.
  • Considerably extreme maneuvers are evaluated to observe the limits of the aircraft by incorporating the coupling and nonlinear effects, contrary to what has been done in the previous works. The efficacy of the proposed method is demonstrated based on these harsh maneuvers.
  • Significantly more ample stable maneuverable volumes are acquired as a result of the comparison between the conventional state limiters and the proposed approach, which is also a pure indicator of the enhancement of the aircraft’s maneuverability without compromising safety.
Consequently, the paramount importance of this study lies in its ability to prevent the loss of control without requiring intensive offline computation besides superiority over the conventional state limiters in terms of maneuverability enhancement. The study’s conclusions demonstrate the significant promise of the proposed method, as it enables the aircraft to maintain stable behavior even subjected to harsh, excessive, and abrupt maneuver requests.
The study is organized as follows: the modeling, employing over-actuated aerodynamic modeling and nonlinear equations of motion for the baseline over-actuated F-16, is briefly presented in Section 2. The flight control architecture is then separated into two distinct but related headlines: the design of the control augmentation system using incremental nonlinear dynamic inversion is given in Section 3.1, and the design of an optimization-based control allocation to handle the over-actuation is presented in Section 3.2. Afterwards, in Section 4, the attainable moment set and its advanced variation are presented, i.e. the incremental attainable moment set. At the core of the study, the architecture for the loss of control prevention, merging Lyapunov’s direct method and the inverse dynamics of the aircraft, is discussed in detail in Section 5. Furthermore, the required subsystems of the architecture are also presented in the same section. The proposed method’s superiority is demonstrated under distinct circumstances and aggressive maneuvers in Section 6. Additionally, a comparison between conventional state limiters and the proposed method is provided. Finally, the concluding remarks and presumable future works are shared in Section 7.
2 Nonlinear Flight Dynamics Model
The baseline aircraft is the F-16 with an over-actuated control suite, meaning that it has five independent control surfaces: two ailerons (left and right), two horizontal tails (left and right), and one rudder. Hereby, the aerodynamic modeling must be adjusted in accordance with this control suite architecture by considering the contributions of each control surface separately. Nevertheless, the over-actuated aerodynamic modeling is adopted from [26, 27]. Each coefficient is expressed as a polynomial function involving the relevant states and control surface deflections. Also, the rigid body 6-degrees-of-freedom nonlinear equations of motion and necessary mass, inertia, actuator, and physical properties are given in [28]. As the cornerstone of this study, the nonlinear flight dynamics model is constructed based on the preceding studies in a high-fidelity manner.
3 Flight Control Architecture
In this section, the design of a control augmentation system using the incremental nonlinear dynamic inversion (INDI) with the fundamental time-scale separation principle is given. Additionally, an optimization-based control allocation design to deal with the over-actuation is discussed.
3.1 Control Augmentation System Design
In this study, the designed control augmentation system has three pilot inputs; angle of attack ($\alpha_{cmd}$), sideslip ($\beta_{cmd}$), and body-axis roll rate ($p_{cmd}$). The fast dynamics and slow dynamics are decomposed in accordance with the time-scale separation. The angle of attack and sideslip are the part of the slow dynamics, whereas the body angular rates are the part of fast dynamics. Therefore, the angle of attack and sideslip are the elements of the outer-loop, and the output of the outer-loop is used as the input of the inner-loop. As a consequence, the dynamics of $\alpha$ and $\beta$ can be such decomposed as the expression given in Eq. (1).
$\begin{bmatrix}\dot{\alpha}\\ \dot{\beta}\\ \end{bmatrix}=\begin{bmatrix}\dot{\alpha}_{0}\\ \dot{\beta}_{0}\\ \end{bmatrix}+G(x_{0},u_{0})\begin{bmatrix}q-q_{0}\\ r-r_{0}\end{bmatrix}$ (1)
If the expression is inverted to derive $q$ and $r$, the following form is obtained as given in Eq. (2).
$\begin{bmatrix}q\\ r\end{bmatrix}=G(x_{0},u_{0})^{-1}\Bigg{(}\begin{bmatrix}\dot{\alpha}\\ \dot{\beta}\\ \end{bmatrix}-\begin{bmatrix}\dot{\alpha}_{0}\\ \dot{\beta}_{0}\\ \end{bmatrix}\Bigg{)}+\begin{bmatrix}q_{0}\\ r_{0}\end{bmatrix}$ (2)
which $G(x_{0},u_{0})$ is given in Eq. (3).
$G(x_{0},u_{0})=\begin{bmatrix}1-\frac{1}{4m\cos\beta}\rho S\bar{c}C_{L_{q}}&-\tan\beta\sin\alpha\\ \frac{\rho S\bar{c}C_{D_{q}}}{4m}\sin\beta(1-\cos\beta)&\frac{\rho SbC_{Y_{r}}}{4m}\cos^{2}\beta-\cos\alpha\\ \end{bmatrix}$ (3)
where $\rho,\bar{c},b,S$ are the density, mean aerodynamic chord length, wing span, and wing area respectively. Also, $C_{L_{q}},C_{D_{q}},C_{Y_{r}}$ are the lift and drag coefficient derivatives with respect to the pitch rate, and side force coefficient derivative with respect to the yaw rate, respectively. Moreover, $[\dot{\alpha}\hskip 4.26773pt\dot{\beta}]^{T}$ is the virtual input and can be obtained using a simple proportional controller as given in Eq. (4).
$\begin{bmatrix}\dot{\alpha}\\ \dot{\beta}\\ \end{bmatrix}=\begin{bmatrix}\omega_{\alpha}(\alpha_{cmd}-\alpha)\\ \omega_{\beta}(\beta_{cmd}-\beta)\\ \end{bmatrix}$ (4)
where $\omega_{\alpha}$ and $\omega_{\beta}$ are the proportional gains for $\alpha$ and $\beta$ channels, respectively. Eventually, the pilot inputs of $\alpha_{cmd}$ and $\beta_{cmd}$ are derived in the form of $q$ and $r$ as the inputs of the inner-loop.
Subsequent to finalizing the outer-loop, the inner-loop is designed based on angular rates; $p$, $q$, and $r$. Furthermore, the outputs of the inner-loop are the control moment coefficients rather than the control surface deflections due to the control allocation. The required moment coefficients can be derived using Euler’s equations of motion as given in Eq. (5).
$\begin{split}\begin{bmatrix}C_{l}\\ C_{m}\\ C_{n}\end{bmatrix}=\frac{J}{\frac{1}{2}\rho V^{2}S}\begin{bmatrix}b&0&0\\ 0&\bar{c}&0\\ 0&0&b\end{bmatrix}^{-1}\Bigg{\{}\begin{bmatrix}\dot{p}\\ \dot{q}\\ \dot{r}\end{bmatrix}+J^{-1}\Bigg{(}\begin{bmatrix}p\\ q\\ r\end{bmatrix}\times J\begin{bmatrix}p\\ q\\ r\end{bmatrix}\Bigg{)}\Bigg{\}}\end{split}$ (5)
which ${[C_{l}\hskip 4.26773ptC_{m}\hskip 4.26773ptC_{n}]^{T}}$ corresponds to the roll, pitch, and yaw moment coefficients as the outputs of the inner-loop. In addition, ${[\dot{p}\hskip 4.26773pt\dot{q}\hskip 4.26773pt\dot{r}]^{T}}$ is the virtual input. As derived previously, the virtual input is given in Eq. (6).
$\begin{bmatrix}\dot{p}\\ \dot{q}\\ \dot{r}\end{bmatrix}=\begin{bmatrix}\omega_{p}(p_{cmd}-p)\\ \omega_{q}(q_{cmd}-q)\\ \omega_{r}(r_{cmd}-r)\\ \end{bmatrix}$ (6)
where $\omega_{p},\omega_{q},\omega_{r}$ are the proportional gains for $p,q$ and $r$ channels, respectively. Note that $q_{cmd}$ and $r_{cmd}$ are derived from the commanded angle of attack and sideslip in the outer loop, whereas $p_{cmd}$ is the direct pilot input. The proportional gains corresponding to each control channel in the proportional controllers are given in Table 1.
Table 1: Proportional controller gains for each control channel
$\omega_{\alpha}$ $\omega_{\beta}$ $\omega_{p}$ $\omega_{q}$ $\omega_{r}$
2.5 1 10 10 5
3.2 Control Allocation Design
The need for a control allocation algorithm due to over-actuation has already been discussed. Thus, to achieve the integrity of the flight control architecture, an optimization-based control allocation is utilized. The objective function represents the minimization of control effort, subject to the position limits. Specifically, the flight control allocation problem can be formulated as given in Eq. (7).
$\min_{u}$ $\frac{1}{2}u^{T}u$ (7)
s.t. $\lVert\Phi(x_{0},u_{0})u-\tau_{c}\rVert=0$
$u_{min}\leq u\leq u_{max}$
where $\Phi(x_{0},u_{0})\in\mathbb{R}^{3\times m}$ is the Jacobian matrix at that instant, where $m$ is the number of control surfaces, and $\tau_{c}$ represents the commanded control moment coefficients. Additionally, the upper and lower bounds, represented by $u_{min}$ and $u_{max}$, significantly influence the derived control surface deflections. Furthermore, the Jacobian matrix is calculated at each time step using the central-difference method to capture the aerodynamic nonlinearity.
The desired control moment coefficient can be achieved by minimizing the control effort with the help of the presented formulation. however, another concern regarding optimization-based control allocation is the selection of a stable, fast-converging, and reliable optimization algorithm. In this study, the selected optimization algorithm is sequential-quadratic programming due to its ability to solve highly complex optimization problems with high reliability, in addition to its stable convergence attitude[29, 30]. Subsequent to finalizing the discussion of the control allocation design, for the sake of clarity, a flowchart representing the entire flight control architecture is illustrated in Fig. 1.
Figure 1: Entire flight control architecture.
As illustrated in Fig. 1, the pilot inputs $\alpha_{cmd}$, $\beta_{cmd}$, and $p_{cmd}$ are subjected to the incremental nonlinear dynamic inversion to derive the control moment coefficients. Afterwards, the control moment coefficient are transferred to the control allocation module to derive the corresponding control surface deflections while minimizing the control effort. All in all, the control surface deflections are injected into the nonlinear flight dynamics module for deriving new states, $x$. Notice that, since the inner-loop is also derived using incremental nonlinear dynamic inversion, the finalized control surface deflections are also in the incremental form. This implies that the control allocation is only responsible for distributing the necessary control moment coefficients on the control surfaces in an optimal manner.
4 Incremental Attainable Moment Set
The attainable moment set (AMS) of an aircraft is a geometrical representation (topologically defines a convex-hull) of the aerodynamic moments that can be generated by the aircraft’s control effectors at various flight conditions. Furthermore, an attainable moment set of an aircraft helps understanding the instantaneous control authority of the aircraft, i.e. controllability of the aircraft [31, 32]. Due to this benefit, the AMS will be used in this study to detect the possibility of loss of control. However, constructing the attainable moment set regarding only the actuator position limits implies that one assesses the attainable moment set of an aircraft, under any arbitrary circumstance, can observe the global achievable control moments. Instead, an alternative path is proposed to give a better insight about the instantaneous controllability during an agile maneuver, i.e. incremental attainable moment set.
The concept and utilization of the incremental attainable moment set (IAMS) is the same as the attainable moment set with a distinction: the boundary formulation. Not only the actuator deflection limits but also the actuator rate limits should be considered to understand whether the aircraft is capable of providing necessary control moment at that time step or not. Therefore, with a slight notion shift, the attainable moment set is reformulated to construct the incremental attainable moment set $\Delta\mathbb{A_{M}}$ as given in Eq. (8).
$\Delta\mathbb{A_{M}}=\{\Delta\tau|\Delta\tau=\Phi\Delta u,\hskip 4.26773pt\underline{\Delta u}\leq\Delta u\leq\overline{\Delta u}\}$ (8)
where $\Delta\tau\in\mathbb{R}^{3}$ is the incremental control moment coefficient matrix, whereas $\Delta u\in\mathbb{R}^{m}$ is the incremental control input. Furthermore, $\underline{\Delta u}$ and $\overline{\Delta u}$ denote the lower and upper boundaries of the control inputs defined as given in Eq. (9).
$\begin{split}\overline{\Delta u}&=\textrm{min}(\dot{u}_{max}\Delta t_{s},u_{max}-u_{0})\\ \underline{\Delta u}&=\textrm{max}(-\dot{u}_{max}\Delta t_{s},u_{min}-u_{0})\end{split}$ (9)
where $\dot{u}_{max}$, $u_{max},u_{min},u_{0}$ are the rate limit, maximum, minimum, and current positions of the related actuator respectively while $\Delta t_{s}$ is the sampling time. Eventually, using the incremental form of the control input allows to rearrange the boundaries of the problem by including the actuator rate limit; moreover, it becomes feasible to obtain more realistic attainable moment set that allows to assess the instantaneous controllability in accordance with the actuator capability.
5 Loss of Control Prevention
The loss of control notion has been discussed in the context of controllability, and this study specifically addresses the prevention of departure. Furthermore, the prevention architecture is divided into two stages: detection and prevention, both of which are proposed in an online manner.
  • Detection: The IAMS serves as the detector for the initiation of the loss of control, indicating the available control authority. The flight control architecture derives the desired control moment vector, which is then compared to the instantaneous IAMS to determine whether the demand falls within the bounds of the available control authority. However, as a safety margin, the acquired instantaneous IAMS is shrunk by a factor of 30%, meaning that only 70% of the available control authority is allowed to be used. This is not a strict requirement but can be modified based on expectations (the greater the shrinkage factor, the more restrictive it is). The specified level is regarded as sufficiently satisfactory in this study, considering the trade-off between maneuverability and loss of control.
  • Prevention: A dynamic (online) command saturation approach is developed in order to prevent the loss of control (i.e. departure) due to the excessive or abrupt pilot inputs. In case of the detection of the violation of the control authority, the dynamic command saturation module is activated and the states of $\alpha$, $p$, $q$ and $r$ are restricted. The dynamic command saturation is designed based on the Lyapunov’s direct method, which will be discussed in the proceeding section.
Consequently, the primary objective is to prevent departure during extreme maneuvers by saturating the pilot commands in the event of a loss of control detection. In this way, executing the most agile maneuvers without entering into an upset condition is desired.
5.1 Lyapunov-based Dynamic Command Saturation Design
The Lyapunov stability theorem has been widely used in the control system design to ensure that a system remains stable in the face of disturbances or uncertainties, and to design control strategies that can stabilize an unstable system. In this section, the derivation of the command saturation will be discussed. Principally, set such a Lyapunov candidate as given in Eq. (10).
$V(\omega)=\omega^{T}J\omega$ (10)
where $\omega=[p\hskip 4.26773ptq\hskip 4.26773ptr]^{T}$ is the angular velocity vector and $J$ is the inertia tensor. Actually, the selected Lyapunov function is an energy-like function, i.e. rotational kinetic energy, and $V(\omega)>0,\forall\omega\in\mathbb{R}^{n}$ since $J\in\mathbb{R}^{3\times 3}$ is a symmetric positive-definite matrix. Moreover, if $\omega^{*}=[0\hskip 4.26773pt0\hskip 4.26773pt0]^{T}$ is the equilibrium point, $V(\omega^{*})=0$. If the time derivative of the Lyapunov candidate is taken as given in Eq. (11),
$\dot{V}(\omega)=2\omega^{T}J\dot{\omega}$ (11)
can be obtained. Recall the Euler’s equations of motion in a compact form as given in Eq. (12).
$J\dot{\omega}=-[\omega\times J\omega]+M$ (12)
If Eq. (12) is plugged in Eq. (11), the following form is obtained as given in Eq. (13).
$\dot{V}(\omega)=2\omega^{T}J\dot{\omega}=2\omega^{T}(-[\omega\times J\omega]+M)$ (13)
The term $2\omega^{T}M$ must be negative, as the criterion $\dot{V}(\omega)<0$ is required for asymptotic stability. Thus, $\dot{V}$ can be transformed to $2\omega^{T}M$, and a moment design specified in Eq. (14) is needed to achieve asymptotic stability in the system.
$M=-K\omega$ (14)
where $K\in\mathbb{R}^{3\times 3}$ is symmetric positive-definite weight matrix. If Eq. (14) is plugged in Eq. (13), the following is obtained.
$\dot{V}(\omega)=-2\omega^{T}K\omega<0$ (15)
Consequently, the stabilizing control moments is derived using Lyapunov’s direct method, and these regulator moments will be utilized to derive the pilot command saturation expressions. By using Euler’s equations of motion, the angular rate saturation can be established through the following relation given in Eq. (16).
$\underbrace{\begin{bmatrix}L\\ M\\ N\end{bmatrix}}_{\begin{subarray}{c}-K\begin{bmatrix}p\\ q\\ r\end{bmatrix}\end{subarray}}=J\underbrace{\begin{bmatrix}\dot{p}\\ \dot{q}\\ \dot{r}\end{bmatrix}_{sat}}_{\begin{subarray}{c}\begin{bmatrix}\omega_{p}(p_{sat}-p)\\ \omega_{q}(q_{sat}-q)\\ \omega_{r}(r_{sat}-r)\end{bmatrix}\end{subarray}}+\begin{bmatrix}p\\ q\\ r\end{bmatrix}\times J\begin{bmatrix}p\\ q\\ r\end{bmatrix}$ (16)
Obviously, the stabilizing moments can be converted into the angular rates as given in the Eq. (17).
$J^{-1}\Bigg{\{}-K\begin{bmatrix}p\\ q\\ r\end{bmatrix}-\begin{bmatrix}p\\ q\\ r\end{bmatrix}\times J\begin{bmatrix}p\\ q\\ r\end{bmatrix}\Bigg{\}}=\begin{bmatrix}\omega_{p}(p_{sat}-p)\\ \omega_{q}(q_{sat}-q)\\ \omega_{r}(r_{sat}-r)\end{bmatrix}$ (17)
Eventually, if $[p\hskip 4.26773ptq\hskip 4.26773ptr]^{T}_{sat}$ is decomposed from the expression given in Eq. (17), the following Eq. (18) can be derived.
$\begin{bmatrix}p_{sat}\\ q_{sat}\\ r_{sat}\end{bmatrix}=\Bigg{\{}J^{-1}\Bigg{\{}-K\begin{bmatrix}p\\ q\\ r\end{bmatrix}-\begin{bmatrix}p\\ q\\ r\end{bmatrix}\times J\begin{bmatrix}p\\ q\\ r\end{bmatrix}\Bigg{\}}\Bigg{\}}\oslash\begin{bmatrix}\omega_{p}\\ \omega_{q}\\ \omega_{r}\end{bmatrix}+\begin{bmatrix}p\\ q\\ r\end{bmatrix}$ (18)
[…]
$\underbrace{\begin{bmatrix}\dot{\alpha}\\ \dot{\beta}\\ \end{bmatrix}}_{\begin{subarray}{c}\begin{bmatrix}\omega_{\alpha}(\alpha_{sat}-\alpha)\\ \omega_{\beta}(\beta_{sat}-\beta)\\ \end{bmatrix}\end{subarray}}=\begin{bmatrix}\dot{\alpha}_{0}\\ \dot{\beta}_{0}\\ \end{bmatrix}+G(x_{0},u_{0})\begin{bmatrix}q-q_{0}\\ r-r_{0}\end{bmatrix}$ (19)
As a result, if $[\alpha\hskip 4.26773pt\beta]^{T}_{sat}$ is decomposed from the expression given in Eq. (19), the following Eq. (20) is derived.
$\begin{bmatrix}\alpha_{sat}\\ \beta_{sat}\\ \end{bmatrix}=\Bigg{\{}\begin{bmatrix}\dot{\alpha}_{0}\\ \dot{\beta}_{0}\\ \end{bmatrix}+G(x_{0},u_{0})\begin{bmatrix}q-q_{0}\\ r-r_{0}\end{bmatrix}\Bigg{\}}\oslash\begin{bmatrix}\omega_{\alpha}\\ \omega_{\beta}\\ \end{bmatrix}+\begin{bmatrix}\alpha\\ \beta\\ \end{bmatrix}$ (20)
In order to make $[\alpha\hskip 4.26773pt\beta]^{T}_{sat}$ stabilizing, the expression of $[q\hskip 4.26773ptr]^{T}$ should be converted into the expression of $[q\hskip 4.26773ptr]_{sat}^{T}$.
$\begin{bmatrix}\alpha_{sat}\\ \beta_{sat}\\ \end{bmatrix}=\Bigg{\{}\begin{bmatrix}\dot{\alpha}_{0}\\ \dot{\beta}_{0}\\ \end{bmatrix}+G(x_{0},u_{0})\begin{bmatrix}q_{sat}-q_{0}\\ r_{sat}-r_{0}\end{bmatrix}\Bigg{\}}\oslash\begin{bmatrix}\omega_{\alpha}\\ \omega_{\beta}\\ \end{bmatrix}+\begin{bmatrix}\alpha\\ \beta\\ \end{bmatrix}$ (21)
where the expression of $[q\hskip 4.26773ptr]_{sat}^{T}$ is derived in Eq.(18). However, only the angle of attack is saturated since the sideslip angle is desired to be zero for all flight circumstances. Consequently, through inductive reasoning, the expression for dynamic command saturation can be derived. By designing the stabilizing control moments, a correlation between them and the dynamic command saturation expressions can be established. Therefore, an online solution for the command saturation is proposed. The only necessity for implementing the proposed online architecture is accurate measurement and a high-fidelity, broad aerodynamic database. The remaining steps involve simple algebraic calculations, so the computational cost depends only on deriving the necessary aerodynamic coefficients. After all, for the sake of clarity, a flowchart of the proposed logic is presented in Fig. 2.
Figure 2: Overview of the proposed loss of control prevention logic
6 Results
[…]
6.1 Abrupt & Agile Maneuver Evaluation
The preliminary assessments of the proposed architecture are presented with two distinct agile maneuvers [^5] For the simulation videos, please visit: https://youtu.be/I524wsm6EPk. The agility metrics of the commands are defined as follows: high amplitude control commands, abrupt application, and simultaneous application. Based on these metrics, angle of attack and roll rate inputs are applied abruptly, with high amplitude, and simultaneously. Furthermore, the static stability degradation has a significant impact on the loss of control; that is, the worst-case scenario is selected, and the center of gravity is located such that the aircraft is statically unstable (at %$\%35\bar{c}$). Subsequently, the effects of the proposed architecture are presented after demonstrating that the aircraft is not capable of performing the desired maneuver without the loss of control prevention architecture. The maneuver properties are detailed in Table 2.
Table 2: Maneuver properties
[…]
Mach Altitude, m
1 0.8 2000 %$\%35\bar{c}$
2 1 500 %$\%35\bar{c}$
The proposed method is examined under various initial conditions, i.e. different dynamic pressure and control authority levels in order to understand the sensitivity of the method’s performance to the alteration of the initial conditions.
6.1.1 Maneuver #1 Examination
Initially, the maneuver performance of the aircraft without the loss of control prevention architecture is depicted in Fig. 3.
Figure 3: State trajectory of the maneuver #1 without loss of control prevention.
Obviously, the aircraft is incapable of executing the demanded maneuver due to the violation of the controllable space, i.e., the attainable moment set, at the instant of input reversals. The surpassing of controllability boundaries initiates a departure from stable flight, resulting in the occurrence of loss of control. Subsequent to stimulation of the loss of control, the control deflections grow irrationally and even reach saturation. Additionally, the discrepancy between the required and achieved control moment coefficients progressively worsens until the simulation is terminated due to numerical issues arises from the loss of control. However, the effects of the loss of control prevention architecture implementation are shown in Fig. 4.
Figure 4: State trajectory of the maneuver #1 with loss of control prevention: indicates the intervention of the Lyapunov assessment.
Furthermore, the dynamic variations in the saturation of the corresponding states of the aircraft are demonstrated in Fig. 5.
Figure 5: Command saturation trajectory of the maneuver #1 with loss of control prevention: indicates the intervention of the Lyapunov assessment.
The attainable moment set history for both loss of control prevention is on and off is demonstrated in Fig. 6.
(a) Without loss of control prevention
(b) With loss of control prevention
Figure 6: Attainable moment set history: (a) without loss of control prevention, (b) with loss of control prevention during maneuver #1.
Attainable moment set illustrations of the maneuver #1 with loss of control prevention indicates the arbitrary instantaneous controllability checks as stated previously, and it is observed that at $t_{3}$ the desired control moment vector violates the attainable moment set, i.e. red, while there is not loss of control prevention. Remember from Figure 3 that the violation stimulates the loss of control and the aircraft departures. However, in case of activation of the loss of control prevention architecture, the demanded control moment coefficient vector is still bounded by the attainable moment set; therefore, any loss of control is avoided. Consequently, the history of the control deflections with demanded moment satisfaction parameter and the 3D trajectory of the maneuver are depicted in Fig. 7 and Fig. 8 respectively.
Figure 7: Control deflections with demanded moment satisfaction parameter of the maneuver #1 with loss of control prevention.
Figure 8: 3D trajectory of the maneuver #1 with loss of control prevention.
Obviously, without the loss of control prevention architecture, the aircraft is incapable of executing the desired maneuver. Eventually, the simulation demonstrates that the aircraft outperforms this maneuver compared to the simulation without the loss of control prevention. Additionally, it is worth highlighting that the maneuver is so agile that the load factor approaches the limit of 10g.
6.1.2 Maneuver #2 Examination
Initially, Fig. 9 depicts the maneuver performance of the aircraft without the loss of control prevention architecture.
Figure 9: State trajectory of the maneuver #2 without loss of control prevention.
Similar to the previous instance, the aircraft is incapable of executing the demanded maneuver due to the violation of the controllable space. Once again, surpassing the controllability boundaries initiates a departure from stable flight, resulting in the occurrence of loss of control. The effects of the loss of control prevention architecture implementation are depicted in Fig. 10.
Figure 10: State trajectory of the maneuver #2 with loss of control prevention: indicates the intervention of the Lyapunov assessment.
Furthermore, the dynamic variations in the saturation of the corresponding states of the aircraft are demonstrated in Fig. 11.
Figure 11: Command saturation trajectory of the maneuver #2 with loss of control prevention.
The attainable moment set histories are demonstrated in Fig. 12.
(a) Without loss of control prevention
(b) With loss of control prevention
Figure 12: Attainable moment set history: (a) without loss of control prevention, (b) with loss of control prevention during maneuver #2.
Consequently, the history of the control deflections with demanded moment satisfaction parameter and the 3D trajectory of the maneuver are depicted in Fig. 13 and Fig. 14 respectively.
Figure 13: Control deflections with demanded moment satisfaction parameter of the maneuver #2 with loss of control prevention.
Figure 14: 3D trajectory of the maneuver #2 with loss of control prevention.
Ultimately, without the loss of control prevention architecture, the aircraft is unable to execute the desired maneuver. It is proven that the aircraft outperforms this maneuver compared to the simulation without the loss of control prevention. It is also noteworthy that the maneuver is so agile that the load factor approaches the limit of 15g.
6.2 Comparison Study: Scheduled State Limiter versus Lyapunov-based Command Saturation
[…]
Figure 15: The comparison for sea level altitude with various Mach number: the volume of the yellow shaded polygon is 0.9504uc whereas the volume of the red shaded polygon is 0.7111uc.
Figure 16: The comparison for 10000ft altitude with various Mach number: the volume of the yellow shaded polygon is 0.7907uc whereas the volume of the red shaded polygon is 0.5094uc.
The most eye-catching distinction of these two separate assessment’s properties is the reduction in the most agile and stable maneuver capability, which is highly related to the control authority alteration due to the available dynamic pressure. Furthermore, an observable distinction in achievable positive and negative roll rates exists because of the asymmetric properties of the aircraft in terms of inertia and engine.
However, perhaps the most significant interpretation of these assessments is that the Lyapunov-based command saturation approach allows for more agile maneuvers without compromising stable flight, which has been the main objective of the study. In terms of expansion, the Lyapunov-based command saturation approach increased the generated volume compared to the conventional approach by 33.66% for the sea level altitude assessment and by 55.21% for the 10000ft altitude assessment. This serves as concrete evidence that more agile and stable flight can be achieved using the proposed method. However, it is noteworthy that, even using the proposed method, not all combinations of angle of attack and roll rate maneuvers can be performed without losing control.
7 Conclusion & Future Works
In this study, an online loss of control prevention methodology is presented for over-actuated F-16 model through merging the Lyapunov’s direct method and inverse dynamics of the aircraft. The flight control architecture is established based on the design of flight control law using incremental nonlinear dynamic inversion and an optimization-based control allocation due to the over-actuation, whereas the pilot inputs are selected as the angle of attack ($\alpha$), sideslip angle ($\beta$), and roll rate ($p$). However, the challenge of the proposed methodology is not only to control the aircraft but also to ensure the flight safety by eliminating the destabilizing effects which has been discussed.
To address this critical issue, a structure with two components is developed: a controllability investigation module (using incremental attainable moment set) and a dynamic command saturation module (merging Lyapunov’s direct method and inverse dynamics of the aircraft). Through the proposed structure, the control commands are dynamically saturated during flight to prevent loss of control in the event of a detection of a loss of control possibility. Prior to applying dynamic saturation, a controllability assessment is conducted using an incremental attainable moment set, which considers actuator limitations such as position and rate saturation. This assessment determines whether the demanded control moment vector falls within the attainable set or not, whereas the allowable control authority is set at 70%, with a 30% safety margin. Violating this set indicates an excess of control authority, and in such cases, the Lyapunov-based dynamic command saturation is activated.
The effectiveness and performance of the proposed architecture are demonstrated through two distinct and challenging maneuvers. It is shown that the aircraft is unable to perform these maneuvers without the proposed loss of control prevention structure. However, by activating the loss of control prevention approach, stable maneuvers can be maintained. Furthermore, a comparison is made between the conventional state limiters and dynamic command saturation, revealing a significant expansion of the stable maneuverable volume using the proposed method (33.66% at the sea level and 55.21% at 10000ft). In conclusion, the proposed architecture extends the safe maneuvering envelope of the aircraft, enabling the execution of most abrupt and agile maneuvers without departing from stable flight.
Further studies are planned to examine the effects of the scheduled gain on the expansion of the stable maneuverable volume using an adaptive Lyapunov gain. Additionally, another aim is to explore the compatibility of the presented method with different baselines, such as commercial aviation or quadcopters. Hence, the plan includes working on the real-time implementation of various baselines and investigating scenarios involving a fault occurrence, utilizing either the presented methodology or an improved version, as a part of future work.